OpenStudy (anonymous):

Solve for t. Clear fractions first: 2/3t+4=1/6t

OpenStudy (anonymous):

What's your first guess?

OpenStudy (anonymous):

Do you divide 2/3 by both sides?

OpenStudy (anonymous):

Nah, you wanna get all the \(t\) terms on one side.

OpenStudy (anonymous):

Do you know how to add/subtract like terms?

OpenStudy (anonymous):

\[t=-\frac{ 1 }{ 8 }\] Right?

OpenStudy (anonymous):

I cannot remember. It has been a long time since I have done algebra..

OpenStudy (anonymous):

Can you do: \[ \frac{1}{6}t-\frac{2}{3}t \]

OpenStudy (anonymous):

\[\frac{ 2 }{ 3t }-\frac{ 1 }{ 6t }=-4\] \[\frac{ 2 }{ 3t }=\frac{ 4 }{ 6t }\] \[\frac{ 4-1 }{ 6t }=-4\]

OpenStudy (precal):

you could just multiply the whole equation (both sides) by 6 (the LCD) to get rid of the fractions

Directrix (directrix):

@LylaJane Clear fractions first : (2/3) t + 4 = (1/6) t You have denominators of 3 and 6. Clear fractions by multiplying each member of the equation by 6. 6*(2/3) t + 6*4 = 6* (1/6) t That yields: 4t + 24 = t. Solve for t.

OpenStudy (yacoub1993):

Can anyone help me Please :) \[-1\frac{ 3 }{ 4 }*\left( -3\frac{ 2 }{ 3 } \right)\]

OpenStudy (precal):

I would use directrix's approach

OpenStudy (texaschic101):

Yacoub1993 .... change mixed numerals to improper fractions, then multiply... you can do this...

OpenStudy (yacoub1993):

ok i made a mistake

OpenStudy (yacoub1993):

its \[-1\frac{ 3 }{ 7 }\times(-3 \frac{ 2 }{ 3 })\]

OpenStudy (texaschic101):

Do you know what - 1 3/7 is when changed to an improper fraction ?

OpenStudy (yacoub1993):

\[-\frac{ 3 }{ 7 }\]

OpenStudy (texaschic101):

no..... To change mixed numbers to improper fractions, you take the whole number and multiply it by the denominator, then add that to the numerator, then put that over the original denominator. Lets do this...1 x 7 = 7 + 3 = 10, then put 10 over the denominator and that makes it - 10/7. This is your improper fraction for -1 3/7. Can you do the other mixed number ?

OpenStudy (yacoub1993):

is it \[-\frac{ 11 }{ 3 }\]

OpenStudy (texaschic101):

yes it is....now just multiply ....- 10/7 x - 11/3 and you will get your answer

OpenStudy (yacoub1993):

i got 110/21 but the choices they gave me doesnt include that

OpenStudy (texaschic101):

you have to turn the improper fraction (110/21)back into a mixed number. Do you know how to do that ?

OpenStudy (yacoub1993):

no u just tought me the opposite

OpenStudy (texaschic101):

Let me show you.... How many times does 21 go into 110 without going over ?

OpenStudy (yacoub1993):

i think 5

OpenStudy (texaschic101):

yes....5 is your whole number. 5 x 21 = 105. Now take the 110 and subtract 105 and what do you get ?

OpenStudy (yacoub1993):

5

OpenStudy (yacoub1993):

hold on

OpenStudy (texaschic101):

yes...that is your numerator. And you put your numerator over the original denominator and you get 5 5/21.

OpenStudy (yacoub1993):

is it \[5\frac{ 5 }{ 21 }\]

OpenStudy (texaschic101):

yes it is ....way to go !!

OpenStudy (yacoub1993):

man dont know how to reward u back

OpenStudy (texaschic101):

you understanding is reward enough :)

OpenStudy (yacoub1993):

ok i have some more question i will post as public help me with that too bro

OpenStudy (texaschic101):

I am not a bro .....lol....I am a chic ...lol :)

OpenStudy (yacoub1993):

sorry sis